OpenStudy (anonymous):

Simplify 15x=15/x^2-1 Isn't my answer 15/(x-1)

OpenStudy (anonymous):

is that an equal sign?

OpenStudy (anonymous):

\[15x=\frac{15}{x^2-1}\]?

OpenStudy (anonymous):

ops I meant"+"

OpenStudy (anonymous):

\[15x+\frac{15}{x^2-1}\] ?

OpenStudy (anonymous):

and your job is to add right?

OpenStudy (anonymous):

I have to simplify the question I believe the answer is 15 over x-1

OpenStudy (anonymous):

not if it is what i wrote above

OpenStudy (anonymous):

\[a+\frac{b}{c}=\frac{ac+b}{c}\]

OpenStudy (anonymous):

ditto on what @satellite73 said. Are you comfortable with using the equation editor? Or can you rewrite it the way you intended really quickly?

OpenStudy (anonymous):

you would need \[\frac{15x(x^2-1)+15}{x^2-1}\] as a first step. do no look to cancel, but rather multiply out in the numerator and combine like terms

OpenStudy (anonymous):

(15)/(x+1)?.. being just to =15

OpenStudy (anonymous):

What @satellite73 wrote above is correct if in fact it we are adding.

OpenStudy (anonymous):

Was the problem written correctly? You can take a picture of the problem from the book and attach an image if needed, otherwise I second satellite.

OpenStudy (anonymous):

Simplify 15x_15 over x^2-1

OpenStudy (anonymous):

Simplify 15x+15 over x^2-1***

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

\[\frac{15x+15}{x^2-1}\] right!!

OpenStudy (anonymous):

got it and your answer is correct \[\frac{15}{x-1}\]

OpenStudy (anonymous):

Yeah you had it right.

OpenStudy (anonymous):

Hahah, Thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I'd try to play around with the equation editor on the bottom left of this message box. It'll help with displaying work cleanly.

OpenStudy (anonymous):

I voted for you @satellite73 :-)

OpenStudy (anonymous):

I have no idea how to use that O.O

OpenStudy (anonymous):

It's cool. It's not mandatory or anything. But if you having any spare time play around with it. It's not as hard as it looks. Take care!

OpenStudy (anonymous):

Thank you for all your help!!

OpenStudy (anonymous):

No sweat!